OpenStudy (anonymous):

Generate the first 5 terms of this sequence: (1 point) f(1) = 2 and f(2) = 3, f(n) = f(n - 1) + f(n - 2), for n > 2. 2, 3, 5, 7, 9 2, 3, 4, 7, 11 2, 3, 5, 9, 11 2, 3, 5, 8, 13

OpenStudy (anonymous):

please help

Nnesha (nnesha):

n = next term(term that u have to find ) so replace n by 3 4 and then 5 \[\huge\rm f(\color{red}{n})=f(\color{red}{n}-1)+(\color{red}{n}-2)\]

Nnesha (nnesha):

to find 3rd term replace n by 3 :)

OpenStudy (anonymous):

 now what do i do?

Nnesha (nnesha):

yep now solve parentheses

OpenStudy (anonymous):

how do i do that?

Nnesha (nnesha):

(3-1) = ? (3-2)= ??

OpenStudy (anonymous):

(3-1)= 2 (3-2)= 1

OpenStudy (anonymous):

what do i do with this though

Nnesha (nnesha):

that's right so we can write \[\huge\rm f(\color{red}{3})=f(\color{red}{2})+f(\color{red}{1})\] right so f(2) = ?? what in the question f(1) = ??

OpenStudy (anonymous):

f(2) = 3 f(1) = 2 right?

Nnesha (nnesha):

yep right so you can replace f(2) by 3 and f(1) by 2 and then add

Nnesha (nnesha):

did you understand ?

OpenStudy (anonymous):

what do you mean by add?

Nnesha (nnesha):

add f(2) + f(1)

OpenStudy (anonymous):

sorry im clueless. would the answer be f(3) then???

Nnesha (nnesha):

f(2) = 3 and f(1) = 2 so f(3) = 3 +2 :-)

Nnesha (nnesha):

 f(2) = 3 so you can replace f(2) by 3 right ?